Four equal amount of belladonna exact containing 1.15%. 1.3%, 1.35% and 1.2% alkaloid, were mixed. What was the percentage strength of the mixture?


Found 2 solutions by Hung, Alan3354:
Answer by Hung(2)

1.15% x 4 = 0.04
1.3% x 4 = 0.002
1.35% x = 0.054
1.2% x 4 = 0.096
0.04+0.002+0.054+0.096=0.398
1.15+1.3+1.35+1.2=5
0.398/5 = 8% This is answer.

Answer by Alan3354(69443)

Four equal amount of belladonna exact containing 1.15%. 1.3%, 1.35% and 1.2% alkaloid, were mixed. What was the percentage strength of the mixture?
-------------
It's the average of the 4.
= (1.15 + 1.3 + 1.35 + 1.2)/4
= 1.25%